# Paul-Erik Haacker, Remco I. Leine, Renu Chaudhary, Kai Diethelm, André Schmidt, Safoura Hashemishahraki

### v1

@@ 3 Liouville-Weyl Differential Equations: A Class of Systems with Periodic Solutions
The example (i) illustrates, that exponentially decaying initial functions are not feasible, as these grow unboundedly in negative time. We will also observe in the next section that decaying solutions of linear versions of (11a) posess an algebraic decay instead of an exponential decay. Exponential growth in case (ii) is indeed possible and will be observed in the two following sections. In example (iii), the initialization function $x_{0}$ is a typical trajectory generated by a linear fractional-order Caputo differential equation with finite lower bound. Although its first derivative is undefined at time t=1t=-1, its forcing term is well defined. Example (iv) is an example showing that a singularity of $x_{0}^{\prime}$ at $t_{0}$ can lead to nonexistence of $\mathcal{F}x_{0}$. This particular example of $x_{0}$ is bounded and uniformly continuous, showing that the space of initial functions of (11a) needs to be chosen differently than in [6]. Example (v) allows for a explicit computation of the forcing term. As this work is motivated by the study of periodic solutions, example (vi) is a periodic initial function. In fact, the Liouville-Weyl differential equation (11a) can possibly posess a periodic solution, in contrast to fractional differential equations with finite lower bound [14]. Indeed, a fractional-order Caputo derivative to finite lower bound of a periodic function is not periodic [20]. A minimal example of a system with a periodic solution is given in the following.
@@
$x(t)=\begin{bmatrix}\sin(t)-\frac{\cos(\alpha\pi/2)}{2\sin(\alpha\pi/2)}\cos(t)\\ \frac{\cos(\alpha\pi/2)}{2\sin(\alpha\pi/2)}\sin(t)+\cos(t)\end{bmatrix}$. (25)
Besides existence of a solution, we show that an initial condition $x_{0}\in X(-\infty,t_{0}]$ generates a trajectory $x\in X(-\infty,t_{1}]$ for some $t_{1}>t_{0}$. The following Lemma yields the properties of continuity and boundedness of $x$, as well as boundedness of $x^{\prime}$ in a neighborhood of $t=t_{1}$.
Lemma 3.8 ([15], Lemma 2.1 and 2.6).
Consider the nonlinear fractional-order Caputo differential equation
${}^{\mathrm{C}}\operatorname{D}_{a}^{\alpha}x_{L}(t)=f_{L}(t,x_{L}(t)),\quad x_{L}(a)=x_{a}\in\mathbb{R}$ (26)
with $\alpha\in(0,1]$, $f_{L}$ continuous and locally Lipschitz w.r.t. its second argument for $t\geq a$. Then, there exists $h>a$ s.t. $x_{L}\in C[a,h]\cap C^{1}(a,h]$.
Proposition 3.9 (Invariance of $X$).
Consider the initial value problem (11) along with its solution $x$ on $(-\infty,t_{1}]$ for some $t_{1}>t_{0}$, i.e. $x(t)$ satisfies the differential equation (11a) for $t\in[t_{0},t_{1}]$ and $x(t)=x_{0}(t)$ for $t\in(-\infty,t_{0}]$ with $x_{0}\in X(-\infty,t_{0}]$. It then holds that $x\in X(-\infty,t_{1}]$.
Proof 3.10.
Lemmata 3.4 and 3.8 directly yield $x\in C(-\infty,t_{1}]\cap L^{\infty}(-\infty,t_{1}]$ and $x^{\prime}\in L^{\infty}[t_{1}-\eta,t_{1}]$ for some $\eta>0$ sufficiently small.
We intend to study stability properties of linear cases of the system (11a) in the subsequent sections. We give the stability definition here, which is equivalent to the one of [6] and the classical literature on functional differential equations [22].
@@ 4 Linear Time-Invariant Fractional-Order Systems
In order to demonstrate the difficulties in finding analytical solutions of fractional-order Caputo differential equations with infinite memory, we first consider linear time-invariant differential equations. Recall the classical initial value problem with zero lower bound [2]
${}^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}z(t)=Az(t),\quad t\geq 0,\quad z(0)=z_{0}\in\mathbb{R}^{n},$ (28)
@@
yields (in the semisimple case) a complete basis of solutions of (28), where $E_{\alpha}$ denotes the one parameter Mittag-Leffler function and $v_{z}\in\mathbb{R}^{n}$. Its Caputo derivative reads $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}z(t)=\lambda z(t)$ [2], yielding the characteristic equation
@@
Solutions $\lambda$ of the characteristic equations (30) are directly related to the stability properties of the trivial solution. This is stated in the following Theorem, where the definitions of stability terms are given in the same reference.
Theorem 4.1 ([2], Theorem 7.20).
@@
where $v_{u}\in\mathbb{R}^{n}$ and $\mathrm{Re}(s)>0$. This is a natural choice, as its Liouville-Weyl derivative reads $^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}u(t)=s^{\alpha}u(t)$ [7], yielding the characteristic equation
@@
Solutions of (31) may be constructed irrespective of the solutions of the characteristic equation (33). As shown in Lemma 3.4, each initial value problem (31) of Liouville-Weyl type may be rewritten as an inhomogenous initial value problem of Caputo type
${}^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}u(t)=Au(t)-\mathcal{F}u_{0}(t),\quad t\geq 0,\quad u(0)=u_{0}(0)$ (37)
with the forcing term $\mathcal{F}u_{0}$ of Definition 3.2. Applying the variation-of-constants formula [2] has the solution
@@
where $\tilde{A}=\operatorname{diag}(A,A)$. Clearly, the two components of this system are decoupled, and each of them is equivalent to the original initial value problem, so this formulation might appear unnecessarily complicated. It has the advantage, however, that we can now directly use the methods and results of [3], which hold for planar systems.
@@
It follows that $Q(s)=0$ for some $s\in\mathbb{C}$ if and only if $s^{\alpha}=A$. But now, for any $s\in\mathbb{C}$ we have that $\arg(s)\in(-\pi,\pi]$, and hence $\arg(s^{\alpha})=\alpha\cdot\arg(s)\in(-\alpha\pi,\alpha\pi]$. Since in our case $\alpha\in(0,1)$, we thus conclude that $|\arg(s^{\alpha})|\neq\pi$, and so the equation $s^{\alpha}=A$ cannot have a solution. Thus, the characteristic function does not possess any zeros in $\mathbb{C}$, and so in particular it does not have any zeros in the closed right half of the complex plane. This observation allows us to invoke [3, Theorem 4] which immediately yields the desired result.
@@
Here we use that, since $E_{\alpha,\alpha}(A(\cdot)^{\alpha})$ is monotonically decreasing whenever $A<0$ (this immediately follows from the complete monotonicity of the function $E_{\alpha,\alpha}(-(\cdot))$ shown in [1, § 4.9.2]), it holds that
@@ 5 Fractional-Order Linear Time-Periodic Systems
and $y_{0}\in X$. The coordinate $y(t)$ is introduced to describe the local behavior of a small perturbation from the periodic solution $x_{\mathrm{p}}$. Thus, asymptotic stability and instability of the periodic solution is given whenever the equilibrium $y\equiv 0$ of (51) is asymptotically stable and unstable, respectively.
The choice of the lower bound -$\infty$ in the differential equation (11a) is necessary when we want to study periodic solutions $x_{\mathrm{p}}(t)=x_{\mathrm{p}}(t+T)$ for all $t$ and some $T>0$ of (11). A version of (11) with zero lower bound cannot have a periodic solution [14]. We intend to study the behavior of the solutions of (51). To the best of our knowledge, there are no results on the form of solutions for this problem, i.e. no Floquet theory. Here, we will attempt to fill that gap.
@@
whenever $\operatorname{Re}\lambda\geq 0$ and where $(\cdot)^{\alpha}$ denotes the principal value, $\alpha\in(0,1)$ [7]. One may compute the fractional derivative of the Floquet form (53) by interchanging the order of summation and fractional differentiation, justified by Assumption (A2),
@@
The technical Assumptions (A1) and (A2) on $J$ and $p$, respectively, ensure that the reformulation into the nonlinear eigenvalue problem (73) is valid. Both assumptions are following the approach of [19] on first-order systems. For a given function $J$, Assumption (A1) is easily checked. However, as the function $p$ is a priori unknown and is only approximated via truncated Fourier series in the following, Assumption (A2) remains a technical assumption.
A nonlinear eigenvalue problem such as (73) as a characteristic equation is not uncommon. We have seen in Section 4, that a similar but simpler eigenvalue problem occurs for time-invariant systems. Another example of nonlinear eigenvalue problems as characteristic equations shows up in delay differential equations [27] for both the time invariant and time periodic cases. Therein, the nonlinearity in the eigenvalue problem is an exponential of the eigenvalue $\lambda$.
The constraint $\operatorname{Re}\lambda\geq 0$ is neccessary, since otherwise the Liouville-Weyl derivative does not exist. This shows, that an extension of Floquet form solutions as in [5] to the system (51) is generally not possible. For a certain class of systems with memory terms, results on forms of solutions are available [5]. However, fractionally damped mechanical systems [6]
@@
just barely escape the requirements of [5], since the memory kernel $K(t,\tau)=(t-\tau)^{-\alpha}/\Gamma(1-\alpha)$ does not satisfy $\int_{-\infty}^{t}|K(t,\tau)|\mathrm{d}\tau<\infty$ for any $t$.
@@
By the structure of the fractional Hill matrix (55), given a solution $\lambda$ which makes $\mathbf{H}_{\infty}^{\alpha}(\lambda)$ rank deficient, $\lambda+ik\omega$ for all $k\in\mathbb{Z}$ is also a solution. Hence, the Floquet exponents come in groups, just as in the integer-order case $\alpha=1$ [4]. This comes as no surprise, as the property is a consequence of the ansatz (53).
@@
${}^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}\tilde{y}(t)$ $=J(t)\tilde{y}(t),\quad t>0$ (69a)
@@
where $J$ is $T$-periodic. Inspired by the form of solutions of LTI first-order DEs and LTI-FODEs with lower bound zero, taking an Ansatz as (53) but with Mittag-Leffler functions, will greatly complicate setting up a Hill-type problem due to the lack of a multiplicative identity of the Mittag-Leffler function [2]
@@
belongs to the class of Nonlinear Eigenvalue Problems (NEPs) described by [8]
@@
where the eigenvalue $z$ may appear nonlinearly, but the corresponding eigenvector $v$ linearly in the equation. Numerical treatment of these problems is extensively discussed in [8].
@@
In the case $\alpha=1$, $\lambda$ is a linear eigenvalue of the Hill matrix. Indeed, $\mathbf{H}_{N}^{(1)}(\lambda)=\mathbf{H}_{N}-\lambda I$ where $\mathbf{H}_{N}=\mathbf{H}_{N}^{(1)}(0)$ is the classical Hill-matrix [4].
@@
It holds that
@@
For the sake of simplicity, consider the scalar case $n=1$. We apply [8, Thm 2.19] noting that the fractional Hill matrix
@@
Figure 3: Numerical solutions of Example 5.12 with constant initial condition $y_{0}(t)=1$ for $t\leq 0$, $\alpha=1/2$, a=1a=-1 and varying parameter $b$. We used Garappa’s implementation fde12.m in Matlab of the FracPECE scheme [9].
@@
(a) $b=1$, converging numerical solution $y(t)$ of (81).
(b) $b=2.5$, diverging numerical solution $y(t)$ of (81).
@@
Figure 6: Results of numerical time integration $y_{\mathrm{sim}}(t)$ of the system (81) with $\alpha=1/2$, a=1a=-1, $N=10$ and $b=2.2$, compared with the solution obtained by the Hill method $y_{\mathrm{Hill}}(t)$. We used the numerical scheme of [10] implemented by Afshin Farhadi for problems of the form $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x=f(t,x)$ to initial condition $x(0)=x_{0}$. We implemented the infinite-memory initial condition $y_{0}$ as a forcing term on the right-hand side.

### v2

@@ 3 Liouville-Weyl Differential Equations: A Class of Systems with Periodic Solutions
The example (i) illustrates that exponentially decaying initial functions are not feasible, as these grow unboundedly in negative time. We will also observe in the next section that decaying solutions of linear versions of (12a) possess an algebraic decay instead of an exponential decay. Exponential growth in case (ii) is indeed possible and will be observed in the two following sections. In example (iii), the initialization function $x_{0}$ is a typical trajectory generated by a linear fractional-order Caputo differential equation with finite lower bound. Although its first derivative is undefined at time t=1t=-1, its forcing term is well defined. Example (iv) is an example showing that a singularity of $x_{0}^{\prime}$ at $t_{0}$ can lead to nonexistence of $\mathcal{F}x_{0}$. This particular example of $x_{0}$ is bounded and uniformly continuous, showing that the space of initial functions of (12a) needs to be chosen differently than in [Hinze2020]. Example (v) allows for an explicit computation of the forcing term. As this work is motivated by the study of periodic solutions, example (vi) is a periodic initial function. In fact, the Liouville-Weyl differential equation (12a) can possibly posess a periodic solution, in contrast to fractional differential equations with finite lower bound [Tavazoei2009]. Indeed, a fractional-order Caputo derivative with finite lower bound of a periodic function is not periodic [Area2016]. A minimal example of a system with a periodic solution is given in the following.
@@
$x(t)=\begin{bmatrix}\sin(t)-\frac{\cos(t)}{2\tan(\alpha\pi/2)}\\ \frac{\sin(t)}{2\tan(\alpha\pi/2)}+\cos(t)\end{bmatrix}$. (28)
We intend to study stability properties of linear cases of the system (12a) in the subsequent sections. We give the stability definition here, which is equivalent to the one of [Hinze2020] and the classical literature on functional differential equations [Hale1977].
@@ 4 Linear Time-Invariant Fractional-Order Systems
In order to demonstrate the difficulties in finding analytical solutions of fractional-order Caputo differential equations with infinite memory, we first consider linear time-invariant differential equations. Recall the classical initial value problem with zero lower bound [Diethelm2010]
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast 0}^{\alpha}}z(t)=Az(t),\quad t\geq 0,\quad z(0)=z_{0}\in\mathbb{R}^{n},$ (30)
@@
yields (in the semisimple case) a complete basis of solutions of (30), where $E_{\alpha}$ denotes the one parameter Mittag-Leffler function and $v_{z}\in\mathbb{R}^{n}$. Its Caputo derivative reads ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast 0}^{\alpha}}z(t)=\lambda z(t)$ [Diethelm2010], yielding the characteristic equation
@@
Solutions $\lambda$ of the characteristic equations (32) are directly related to the stability properties of the trivial solution. This is stated in the following Theorem taken from [Diethelm2010, Theorem 7.20] where the definitions of stability terms are given in the same reference.
Theorem 4.1.
@@
where $v_{u}\in\mathbb{R}^{n}$ and $\mathrm{Re}(s)>0$. This is a natural choice, as its Liouville-Weyl derivative reads $^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}u(t)=s^{\alpha}u(t)$ [samko1993fractional], yielding the characteristic equation
@@
Solutions of (33) may be constructed irrespective of the solutions of the characteristic equation (35). As discussed in Proposition 3.5, each initial value problem (33) of Liouville-Weyl type may be rewritten as an inhomogenous initial value problem of $\operatorname{D}_{\ast 0}^{\alpha}$-type
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast 0}^{\alpha}}u(t)=Au(t)-\mathcal{F}u_{0}(t),\quad t\geq 0,\quad u(0)=u_{0}(0)$ (39)
with the forcing term $\mathcal{F}u_{0}$ of Definition 3.2. Applying the variation-of-constants formula [Diethelm2010], problem (39) has the solution
@@
where $\tilde{A}=\operatorname{diag}(A,A)$. Clearly, the two components of this system are decoupled, and each of them is equivalent to the original initial value problem, so this formulation might appear unnecessarily complicated. It has the advantage, however, that we can now directly use the methods and results of [Diethelm22], which hold for planar systems.
@@
It follows that $Q(s)=0$ for some $s\in\mathbb{C}$ if and only if $s^{\alpha}=A$. But now, for any $s\in\mathbb{C}$ we have that $\arg(s)\in(-\pi,\pi]$, and hence $\arg(s^{\alpha})=\alpha\cdot\arg(s)\in(-\alpha\pi,\alpha\pi]$. Since in our case $\alpha\in(0,1)$, we thus conclude that $|\arg(s^{\alpha})|\neq\pi$, and so the equation $s^{\alpha}=A$ cannot have a solution. Thus, the characteristic function does not possess any zeros in $\mathbb{C}$, and so in particular it does not have any zeros in the closed right half of the complex plane. This observation allows us to invoke [Diethelm22, Theorem 4] which immediately yields the desired result.
@@
Here we use that, since $E_{\alpha,\alpha}(A(\cdot)^{\alpha})$ is monotonically decreasing whenever $A<0$ (this immediately follows from the complete monotonicity of the function $E_{\alpha,\alpha}(-(\cdot))$ shown in [GorenfloEtAl2020, § 4.9.2]), it holds that
@@ 5 Fractional-Order Linear Time-Periodic Systems
and $y_{0}\in{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X(-\infty,t_{0}]}$. The coordinate $y(t)$ is introduced to describe the local behavior of a small perturbation from the periodic solution $x_{\mathrm{p}}$. Thus, asymptotic stability and instability of the periodic solution is given whenever the equilibrium $y\equiv 0$ of (53) is asymptotically stable and unstable, respectively.
The choice of the lower bound -$\infty$ in the differential equation (12a) is necessary when we want to study periodic solutions $x_{\mathrm{p}}(t)=x_{\mathrm{p}}(t+T)$ for all $t$ and some $T>0$ of (12). A version of (12) with zero lower bound cannot have a periodic solution [Tavazoei2009].
We intend to study the behavior of the solutions of (53). To the best of our knowledge, there are no results on the form of solutions for this problem, i.e. no Floquet theory. Here, we will attempt to fill that gap.
@@
whenever $\operatorname{Re}\lambda\geq 0$ and where $(\cdot)^{\alpha}$ denotes the principal value, $\alpha\in(0,1)$ [samko1993fractional]. One may compute the fractional derivative of the Floquet form (55) by interchanging the order of summation and fractional differentiation, justified by Assumption (A2),
@@
The technical Assumptions (A1) and (A2) on $J$ and $p$, respectively, ensure that the reformulation into the nonlinear eigenvalue problem (75) is valid. Both assumptions are following the approach of [Zhou2002] on first-order systems. For a given function $J$, Assumption (A1) is easily checked. However, as the function $p$ is a priori unknown and is only approximated via truncated Fourier series in the following, Assumption (A2) remains a technical assumption.
A nonlinear eigenvalue problem such as (75) as a characteristic equation is not uncommon. We have seen in Section 4, that a similar but simpler eigenvalue problem occurs for time-invariant systems. Another example of nonlinear eigenvalue problems as characteristic equations shows up in delay differential equations [Stepan2011] for both the time invariant and time periodic cases. Therein, the nonlinearity in the eigenvalue problem is an exponential of the eigenvalue $\lambda$.
The constraint $\operatorname{Re}\lambda\geq 0$ is neccessary, since otherwise the Liouville-Weyl derivative does not exist. This shows, that an extension of Floquet form solutions as in [Traversa2020] to the system (53) is generally not possible. For a certain class of systems with memory terms, results on forms of solutions are available [Traversa2020]. However, fractionally damped mechanical systems [Hinze2020]
@@
just barely escape the requirements of [Traversa2020], since the memory kernel $K(t,\tau)=(t-\tau)^{-\alpha}/\Gamma(1-\alpha)$ does not satisfy $\int_{-\infty}^{t}|K(t,\tau)|\mathrm{d}\tau<\infty$ for any $t$.
@@
By the structure of the fractional Hill matrix (57), given a solution $\lambda$ which makes $\mathbf{H}_{\infty}^{\alpha}(\lambda)$ rank deficient, $\lambda+ik\omega$ for all $k\in\mathbb{Z}$ is also a solution. Hence, the Floquet exponents come in groups, just as in the integer-order case $\alpha=1$ [Bayer2023a]. This comes as no surprise, as the property is a consequence of the ansatz (55).
@@
$\operatorname{D}_{\ast 0}^{\alpha}\tilde{y}(t)$ $=J(t)\tilde{y}(t),\quad t>0$ (71a)
@@
where $J$ is $T$-periodic. Inspired by the form of solutions of LTI first-order DEs and LTI-FODEs with lower bound zero, taking an Ansatz as (55) but with Mittag-Leffler functions, will greatly complicate setting up a Hill-type problem due to the lack of a multiplicative identity of the Mittag-Leffler function [Diethelm2010]
@@
belongs to the class of Nonlinear Eigenvalue Problems (NEPs) described by [Guettel2017]
@@
where the eigenvalue $z$ may appear nonlinearly, but the corresponding eigenvector $v$ linearly in the equation. Numerical treatment of these problems is extensively discussed in [Guettel2017].
@@
In the case $\alpha=1$, $\lambda$ is a linear eigenvalue of the Hill matrix. Indeed, $\mathbf{H}_{N}^{(1)}(\lambda)=\mathbf{H}_{N}-\lambda I$ where $\mathbf{H}_{N}=\mathbf{H}_{N}^{(1)}(0)$ is the classical Hill-matrix [Bayer2023a].
@@
For a quadratic matrix $M$, it holds that $\operatorname{rank}M=\operatorname{rank}\overline{M}$ and we may take $M=\mathbf{H}_{N}^{\alpha}(\overline{\lambda})$
@@
For the sake of simplicity, consider the scalar case $n=1$. We apply [Guettel2017, Thm 2.19] noting that the fractional Hill matrix
@@
(a) $b=1$.
(b) $b=2$.
Figure 3: Numerical solutions of Example 5.12 with constant initial condition $y_{0}(t)=1$ for $t\leq 0$, $\alpha=1/2$, a=1a=-1 and varying parameter $b$. We used Garappa’s implementation fde12.m in Matlab of the FracPECE scheme [diethelm1998fracpece].
@@
(a) $b=1$.
(b) $b=2.5$.
@@
Figure 6: Results of numerical time integration $y_{\mathrm{sim}}(t)$ of the system (83) with $\alpha=1/2$, a=1a=-1, $N=10$ and $b=2.2$, compared with the solution obtained by the Hill method $y_{\mathrm{Hill}}(t)$. We used the numerical scheme of [Chaudhary2024] implemented by Afshin Farhadi for problems of the form $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x=f(t,x)$ to initial condition $x(0)=x_{0}$. We implemented the infinite-memory initial condition $y_{0}$ as a forcing term on the right-hand side.